OpenStudy (anonymous):

To start off the challenge, the radio station has placed pennies in a cylindrical glass jar. Each penny is 0.75 inches in diameter and 0.061 inches thick. If the cylindrical glass jar containing the pennies has a diameter of 6 inches and a height of 11.5 inches, how many pennies can fit inside the jar? You must show all work to receive credit.

OpenStudy (tylerd):

theres a few ways we can do this, the simplest way is to determine the volume of each penny and divide the volume of the jar by the volume of the penny to = how many pennies can fit.

OpenStudy (tylerd):

of course pennies are not liquid

OpenStudy (tylerd):

so it wouldnt be exactly accurate to say that.

OpenStudy (anonymous):

okay so what should I do

OpenStudy (tylerd):

first calculate the volume of a penny so a penny is a cylinder shape.

OpenStudy (tylerd):

V=pi(radius)^2(height)

OpenStudy (anonymous):

I got 325.15 and 269.49 are those right?

OpenStudy (tylerd):

well the diameter of a penny is 0.75 inches, so the radius would be half that, so 0.375 inches

OpenStudy (tylerd):

the height or (thickness) is 0.061 inches

OpenStudy (anonymous):

were those volume right?? haha

OpenStudy (tylerd):

\[Volume=\pi(0.375)^2(0.061)\]

OpenStudy (tylerd):

so no, the volume of a penny is approximately 0.027 inches cubed

OpenStudy (anonymous):

got 26.93.....

OpenStudy (tylerd):

\[\pi \times 0.140625 \times 0.061\]

OpenStudy (tylerd):

the 0.140625 = 0.375^2

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

what about the next volume I have to find what do I plug in for that

OpenStudy (tylerd):

so the jar is also a cylinder. with diameter 6 and radius is half the diameter so the radius is 3. height is 11.5 inches

OpenStudy (anonymous):

so how would I plug that in the formula

OpenStudy (tylerd):

volume = pi times height times radius squared \[\pi \times 11.5 \times 3^2\]

OpenStudy (anonymous):

so 324.99?

OpenStudy (tylerd):

i get 325.15

OpenStudy (anonymous):

how? I did 3.14x11.5x3x3

OpenStudy (tylerd):

pi is a bit longer then 3.14, i used the alot more decimal places 3.1415926535897932384626433832795

OpenStudy (tylerd):

infact its infinitely longer

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so then what are the two volumes again? because I know u divide the bigger by the smaller one

OpenStudy (tylerd):

yes

OpenStudy (anonymous):

so what were the two volumes??

OpenStudy (tylerd):

0.027 inches cubed and 325.15 inches cubed

OpenStudy (tylerd):

but dont pay attention to the cubed part just divide 325.15/0.027

OpenStudy (anonymous):

okay thank you so much!!! and I have another question that gos with this

OpenStudy (tylerd):

the thing is, you should say this is the answer if we dont factor in the spacing between pennies. that would be alot more work, and your teacher probably isnt expecting you to do all that anyway.

OpenStudy (anonymous):

1.For the Penny Problem, how much empty space should exist inside the jar after being filled to capacity with pennies? Why doesn't this amount of space actually exist in the jar?

OpenStudy (tylerd):

well what we basically said when we did the volumes and divided them is. is you take 12042.59 pennies, and melt them down into a liquid, pour the liquid into the jar and the jar would be full. so in this case, there is no air in the jar, its perfectly filled up.

OpenStudy (anonymous):

okay and one more question

OpenStudy (tylerd):

i think ill leave the rest for you to do, cant just give out all the answers